Consider an object projected straight upward from z=0 with an initial speed of u. Assume that the object is subjected to a linear drag force of the form D=-mKv

a)From the first principle show that t=1/K In(g+Ku/g+Kv) and use this result to find the time taken to reach the maximum height

b)Find the expression for the speed v as a function of time and find the limiting speed


1
Expert's answer
2022-04-01T08:41:45-0400

Answer

a) for these type of D=-mKv and t=1/K In(g+Ku/g+Kv) case force is calculate by using by second law of newton

"F=ma\\
m\frac{dv}{dt}=ma"

b) putting at value of t=1/K In(g+Ku/g+Kv) we get expression for velocity of particle.